The appropriate discount rate for the following cash flows is 9 percent compounded quarterly.
Year Cash Flow
1        $700               
2          600               
3       0               
4 1,400               
What is the present value of the cash flows?
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

700=P1*(1+0.09/4)^(4*1)

P1=700/(1+0.09/4)^(4*1)

=700*0.914843345

=$640.39(Approx).

600=P2*(1+0.09/4)^(4*2)

P2=600/(1+0.09/4)^(4*2)

=600*0.836938346

=$502.16(Approx)

1400=P4*(1+0.09/4)^(4*4)

P4=1400/(1+0.09/4)^(4*4)

=1400*0.700465795

=$980.65(Approx).

Hence total present value=640.39+$502.16+$980.65

=$2123.21(Approx).